OpenStudy (anonymous):

Consider the Quadratic function f(x)=−x2+10x−16. what it's vertex? x-intercepts? y-intercepts?

OpenStudy (anonymous):

vertex always has first coordinate \(-\frac{b}{2a}\)

OpenStudy (anonymous):

in your case you get \(-\frac{10}{2\times(-1)}=5\) second coordinate is what you get when you replace \(x\) by 5

OpenStudy (anonymous):

so x=5 and y=24 ?

OpenStudy (anonymous):

x=5 and y=9 !!

OpenStudy (anonymous):

\[-5^2+10\times 5-16=-25+50-16=25-16=9\]

OpenStudy (anonymous):

so vertex is \((5,9)\)

OpenStudy (anonymous):

x-intercepts?

OpenStudy (anonymous):

set \[-x^2+10x-16=0\] and solve for \(x\) may be easier to solve \[x^2-10x+16=0\]

OpenStudy (anonymous):

try factoring, because this one factors

OpenStudy (anonymous):

ok wbu y-intercept?

OpenStudy (anonymous):

?